### v2

# Efficient Rectified Flow for Image Fusion

Abstract
Image fusion is a fundamental and important task in computer vision, aiming to combine complementary information from different modalities to fuse images. In recent years, diffusion models have made significant developments in the field of image fusion. However, diffusion models often require complex computations and redundant inference time, which reduces the applicability of these methods. To address this issue, we propose RFfusion, an efficient one-step diffusion model for image fusion based on Rectified Flow. We incorporate Rectified Flow into the image fusion task to straighten the sampling path in the diffusion model, achieving one-step sampling without the need for additional training, while still maintaining high-quality fusion results. Furthermore, we propose a task-specific Variational Autoencoder (VAE) architecture tailored for image fusion, where the fusion operation is embedded within the latent space to further reduce computational complexity. To address the inherent discrepancy between conventional reconstruction-oriented VAE objectives and the requirements of image fusion, we introduce a two-stage training strategy. This approach facilitates the effective learning and integration of complementary information from multi-modal source images, thereby enabling the model to retain fine-grained structural details while significantly enhancing inference efficiency. Extensive experiments demonstrate that our method outperforms other state-of-the-art methods in terms of both inference speed and fusion quality. Code is available at https://github.com/zirui0625/RFfusion.
1 Introduction
In computer vision, image fusion is an important task aimed at merging two images from different modalities to obtain a fused image that contains complementary information from both modalities. Image fusion has wide applications across various scenarios. Infrared and visible image fusion (IVIF) liu2022target ; liu2023multi ; xu2020u2fusion ; ma2019fusiongan ; liu2024promptfusion aims to enhance perception under adverse conditions by integrating the detailed information from visible images with the thermal radiation characteristics of infrared images. Medical image fusion (MIF) xu2021emfusion ; li2023gesenet focuses on mitigating the information discrepancies between MRI and CT modalities to provide more comprehensive and accurate diagnostic support. Multi-exposure image fusion (MEF) ma2019deep ; ma2017multi and multi-focus image fusion (MFF) zhang2021deep ; liu2015multi focus on merging images with different exposures and different focal planes, to synthesize high-quality photographic images.
In recent years, with the advent of Denoising Diffusion Probabilistic Models (DDPMs) ho2020denoising , diffusion-based methods have been widely adopted across various computer vision tasks, including image fusion. DDPMs learn the denoising process from noisy observations back to clean images over the data distribution, thereby acquiring the ability to generate high-quality images. Compared to traditional fusion approaches li2023text ; liu2025dcevo , diffusion-based methods zhao2023ddfm ; cao2024conditional ; zhang2024text ; yue2023dif not only effectively integrate information from multiple source images but also significantly enhance the visual quality of the fused results. Benefiting from the powerful priors encoded in pre-trained diffusion models, these approaches demonstrate great potential in multi-task fusion scenarios, where a single unified framework can be adapted to various fusion tasks with remarkable performance.
Figure 1: Efficiency comparison with the state-of-the-art diffusion-based methods.
Although diffusion-based methods have achieved remarkable progress in image fusion tasks, their long inference times pose significant challenges for real-world applications. Recently proposed approaches such as DDFM zhao2023ddfm and CCF cao2024conditional introduce fusion priors into the sampling process of diffusion models, effectively improving fusion quality. However, these methods typically require hundreds of sampling steps to achieve satisfactory results. Reducing the number of steps to improve efficiency often leads to a substantial drop in fusion performance. To accelerate inference in diffusion models, several strategies such as distillation and latent space diffusion have been widely explored. Nonetheless, their application to image fusion remains limited. First, existing distillation methods can enable single-step sampling but usually require fine-tuning tailored to specific model architectures and datasets, lacking generalization across diverse fusion tasks. Second, while latent space diffusion methods based on Variational Autoencoder (VAE) can significantly reduce computational costs, their training objective primarily targets image reconstruction rather than image fusion, leading to considerable challenges when applied directly to fusion scenarios. Therefore, it is imperative to develop a sampling acceleration method that is tailored to image fusion tasks, capable of preserving fusion quality while maintaining generalizability.
To address these challenges, we propose a novel method named RFfusion, which introduces the Rectified Flow mechanism into image fusion tasks for the first time. RFfusion significantly accelerates the inference process of diffusion models without requiring additional training and exhibits strong generalization across multiple fusion tasks. Specifically, we leverage Rectified Flow to construct a linear trajectory between the input images and the target fused image, embedding prior knowledge of the fused image during the sampling process to achieve efficient and high-quality single-step inference. Moreover, we incorporate the sampling process into the latent space and propose a two-stage training strategy to address the objective mismatch between VAE reconstruction and fusion tasks. In the first stage, we fine-tune the VAE to better capture critical features needed for image fusion. In the second stage, the optimized VAE is integrated into the overall fusion framework for joint training, further enhancing the model’s adaptability to fusion scenarios. Extensive experimental results demonstrate that RFfusion not only substantially reduces the number of inference steps and improves computational efficiency, but also outperforms existing state-of-the-art methods across multiple standard image fusion benchmarks. Our contributions can be summarized as:
  • We propose a novel Efficient Rectified Flow image fusion (RFfusion) framework that enables one-step sampling across various fusion tasks without requiring additional training, significantly reducing computational cost and inference time while achieving high-quality fused images.
  • We introduce the image fusion task into the latent space to effectively reduce computational cost. To address the objective discrepancy between the reconstruction-oriented training of VAE and the specific requirements of image fusion, we propose a two-stage training strategy to enhance the VAE’s adaptability to fusion tasks.
  • Extensive experiments demonstrate that our method significantly improves inference speed compared to other diffusion-based approaches. Meanwhile, it also achieves superior fusion performance and shows strong adaptability across various fusion tasks, demonstrating excellent generalization capability..
2 Related works
In this section, we first review influential image fusion algorithms from recent years. Then, we introduce the applications of Rectified Flow in various fields, especially in low-level vision.
Image fusion Image fusion combines images from different modalities to create a single image with complementary information. Traditional methods gu2018recent ; amin2019ensemble ; li2017pixel use convolutional neural networks to achieve image fusion. The Transformer vaswani2017attention ; li2024contourlet , has also advanced the field of image fusion, particularly when combined with CNNs for multi-modal fusion zhao2023cddfuse . Recently, diffusion models have gained attention in low-level vision tasks xia2023diffir ; rombach2022high ; zhao2023ddfm ; cao2024conditional for their strong generative power, also being applied to image fusion tasks. DDFM zhao2023ddfm , using a Denoising Diffusion Probabilistic Model, has shown promising results in infrared-visible and medical image fusion but struggles with adapting to different scenarios. To address this, CCF cao2024conditional proposed a controllable diffusion-based fusion framework, which can optimize the fusion process but still faces challenges like excessive sampling steps.
Rectified flow Liu liu2023flow first proposed the Rectified Flow method, which generates high-quality images by straightening the path between two data distributions, requiring only one or a few sampling steps. InstaFlow liu2023instaflow applies Rectified Flow to text-to-image (T2I) models, using the same approach to straighten the trajectories of probability flows, enabling it to generate high-quality images in a single step. FlowGrad liu2023flowgrad backpropagates gradients along the ODE trajectory, effectively enabling control over the generated content of a pre-trained Rectified Flow model. Recently, some Rectified Flow-based methods li2025one ; zhu2024flowie ; li2025difiisr ; mei2025efficient have also been applied to low-level vision tasks for model acceleration. FlowIE zhu2024flowie constructs a linear many-to-one transport mapping using conditioned Rectified Flow to achieve efficient image enhancement. FluxSR li2025one leverages Rectified Flow to distill diffusion model priors, enabling one-step real-world image super-resolution.
3 Preliminary
Rectified Flow Traditional diffusion models are trained by predicting the noise added during the forward process, enabling the model to generate high-quality images from Gaussian noise. However, this process typically requires multiple sampling steps, which significantly prolongs the inference time. In general, the forward process can be represented as:
$x_{t}=a_{t}x_{0}+b_{t}\epsilon,\quad\epsilon\sim\mathcal{N}(0,1).$ (1)
$a_{t}$ and $b_{t}$ satisfy $a_{t}=1$, $b_{t}=0$ and $a_{t}=0$, $b_{t}=1$. In DDPM, this formula can be expressed as:
$x_{t}=\sqrt{\bar{a}_{t}}x_{0}+\sqrt{1-\bar{a}_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,1).$ (2)
Different from DDPM, Rectified Flow treats the forward process as a transformation between two data distributions, which can be seen as a transformation between Gaussian noise and the real image distribution in this case. The goal of Rectified Flow is to train a model $v_{\theta}$ to predict the velocity $v_{t}(x_{t})$ along the path at step $t$ as
$\mathcal{L}_{\text{RF}}(\theta)=\mathbb{E}_{t,x_{t}}\left\|v_{\theta}(x_{t},t)-v_{t}(x_{t})\right\|^{2}.$ (3)
Rectified Flow views the forward process as a straight path between the real data distribution and the noise distribution, and its noise addition formula can be derived through linear interpolation. According to Equation 1, we can derive:
$x_{t}=(1-t)x_{0}+t\epsilon,\quad\epsilon\sim\mathcal{N}(0,1).$ (4)
In this case, $v_{t}(x_{t})$ can be expressed as:
$x_{t}^{\prime}=v_{t}(x_{t})=\frac{\epsilon-x_{t}}{1-t}=\epsilon-x_{0}.$ (5)
Therefore, according to Equation 3, the training objective of Rectified Flow can be derived as:
$\mathcal{L}_{\text{RF}}(\theta)=\mathbb{E}_{t,x_{t},\epsilon}\left\|v_{\theta}(x_{t},t)-(\epsilon-x_{0})\right\|_{2}^{2}.$ (6)
By training a neural network on a large-scale dataset, the output of the network, $v_{\theta}(x_{t},t)$, is encouraged to closely match the training target $\epsilon-x_{0}$. This enables the model to find the shortest path between two data distributions, significantly accelerating the sampling process.
Variational Autoencoder In diffusion models, Variational Autoencoders (VAEs) are commonly employed to learn low-dimensional latent representations of data. By encoding images into latent spaces, VAEs enable the diffusion process to operate in a more compact latent space, significantly reducing computational costs while improving modeling efficiency. The training objective of a VAE can be formulated as:
$\mathcal{Q}^{*}=\arg\min_{E,G,\mathcal{Z}}\max_{D}\mathbb{E}_{x\sim p(x)}\left[\mathcal{L}_{\text{VQ}}(E,G,\mathcal{Z})+\lambda\mathcal{L}_{\text{GAN}}(\{E,G,\mathcal{Z}\},D)\right],$ (7)
where $E$ and $G$ represent the encoder and decoder, respectively, $\mathcal{Z}$ denotes the discrete codebook, and $D$ is the GAN discriminator. This formulation allows the VAE to produce compact yet semantically meaningful latent codes, which serve as an efficient and expressive latent space for subsequent diffusion modeling.
Figure 2: Illustration of the train and inference pipeline of our methods.
4 Methods
In this section, we first introduce how existing methods utilize diffusion models to achieve image fusion. Then, we describe how we incorporate Rectified Flow into the fusion task. Finally, we present the task-specific Variational Autoencoder (VAE) specifically designed for the fusion task, including the two-stage training strategy of VAE and the loss function used for guidance. The pipeline of our method is illustrated in Figure 2.
4.1 Implementation of Fusion Methods in Diffusion Models
Previous image fusion methods based on diffusion models typically leverage prior knowledge acquired through pre-trained diffusion models to generate high-quality fused images. Inspired by the work presented in chung2023diffusion , these methods incorporate fusion image information into the sampling process via posterior sampling mechanisms of diffusion models, thus effectively guiding the generation of the fused image. During this sampling process, fusion information is progressively integrated and validated, ultimately achieving high-quality image fusion. The specific mathematical formulation can be expressed as follows:
$p_{\theta}\left(f_{(0:T)}\mid i,v\right)=p(f_{T})\prod_{t=1}^{T}p_{\theta}\left(f_{t-1}\mid f_{t},i,v\right),$ (8)
where $f_{0}$ is the fused result and $f_{T}$ is the initial sampling image, usually Gaussian noise. Additionally, the corresponding posterior sampling can be solved using a Stochastic Differential Equation (SDE), and through Bayes’ theorem, we can derive:
$\nabla_{f_{t}}\log p_{t}(f_{t}\mid i,v)=\nabla_{f_{t}}\log p_{t}(f_{t})+\nabla_{f_{t}}\log p_{t}(i,v\mid f_{t}),$ (9)
where $\nabla_{f_{t}}\log p_{t}(f_{t})$ can be obtained via the SDE formulation, inspired by chung2023diffusion , $\nabla_{f_{t}}\log p_{t}(i,v\mid f_{t})$ can be expressed as:
$\nabla_{f_{t}}\log p_{t}(i,v\mid f_{t})\approx\nabla_{f_{t}}\log p_{t}(i,v\mid\tilde{f}_{0\mid t})\approx\rho\nabla_{f_{t}}\|i,v-\mathcal{M}(\hat{f}_{0}(f_{t}))\|_{2}^{2}.$ (10)
Therefore, the fusion prior can be incorporated into the sampling process by computing the observations $\|i,v-\mathcal{M}(\hat{f}_{0}(f_{t}))\|$ between the fused image and the input images. In this manner, the high-quality image generation capability of diffusion models is effectively leveraged to achieve image fusion. A more detailed derivation of the formulas is provided in Appendix A.
4.2 One Step Effective Fusion Network
To enable efficient one-step image fusion, inspired by liu2022flow , we adopt Rectified Flow for the fusion task. Specifically, we utilize a pre-trained model based on Rectified Flow sampling to generate high-quality fused images. Notably, we observe that using visible-light images as input, rather than pure Gaussian noise, leads to improved fusion performance. The sampling process in our method can be formally expressed as:
$f_{t-\Delta t}=f_{t}-\Delta t\times v(f_{t},t)\quad v_{t}(f_{t})=\frac{\epsilon-f_{t}}{1-t}\quad t\in[1,0].$ (11)
Followed by DDFM zhao2023ddfm , we incorporate the inference results of the Expectation-Maximization (EM) algorithm into the sampling process of the diffusion model, thereby injecting the prior of the fused image into the diffusion model through posterior sampling, and achieving image fusion. This process can be formulated as:
$p_{\theta}(f_{0}\mid i,v)=\int p(f_{t})\,\delta\left(f_{0}-\left(f_{t}-\Delta t\cdot v_{\theta}(f_{t}\mid i,v)\right)\right)\,df_{t},$ (12)
where $p(f_{t})$ denotes the initial distribution, while the Dirac delta function $\delta$ ensures that the output $f_{0}$ is strictly determined by the input $f_{t}$ and the velocity field $v_{\theta}(f_{t}\mid i,v)$. It is important to note that Rectified Flow leverages an Ordinary Differential Equation (ODE) framework, meaning that no stochastic noise is injected during the sampling process. Instead, data is deterministically transformed from an initial distribution to the target distribution by optimizing a continuous velocity field $v_{\theta}(f_{t}|i,v)$. Consequently, Equation 9 in our method can be reformulated as
$v_{\theta}(f_{t}|i,v)=v_{\theta}(f_{t})+\nabla_{f_{t}}\log p(i,v\mid f_{t})\approx v_{\theta}(f_{t})+\nabla_{f_{t}}\log p_{t}(i,v\mid\tilde{f}_{0\mid t}).$ (13)
In this way, we transfer Rectified Flow to the image fusion task, achieving an efficient single-step image fusion method without requiring additional training.
4.3 VAE Autoencoder for Image Fusion
LDM rombach2022high was the first to introduce generative diffusion models into the latent space, leveraging the powerful encoding capability of Variational Autoencoder (VAE) to perform image tasks in the latent space, significantly reducing inference costs while achieving high visual fidelity.
Inspired by this work, we introduce VAE into the image fusion task to enable image generation in the latent space. However, applying VAE-based approaches to image fusion faces two key challenges: (1) Previous methods for image reconstruction typically focus on pixel-level visual fidelity, whereas the core of image fusion lies in capturing complementary semantic information across different modalities. (2) Unlike traditional reconstruction tasks where the objective is to recover the original input, image fusion requires decoding a fused image that integrates information from multiple input modalities. Due to inherent differences between the input images and the desired fused output, this discrepancy poses a significant challenge for the direct application of pretrained VAE in image fusion tasks. To address the aforementioned challenges, we propose a two-stage training strategy to effectively adapt VAE architectures for the image fusion task.
VAE training stage I To address Challenge I, we devise a training strategy based on frequency similarity. Specifically, prior research has demonstrated that the complementary semantic information emphasized in image fusion is often closely correlated with the high- and low-frequency components of the input images. Leveraging this insight, we introduce a frequency similarity loss and fine-tune only the VAE encoder and decoder, without involving Rectified Flow sampling or the image fusion process. As a result, the training procedure closely resembles that of conventional image reconstruction. Followed by Equation 7, the corresponding training goal is formulated as follows:
[…]
where $\mathcal{L}_{\text{VQ}}$, $\mathcal{L}_{\text{GAN}}$ are followed by esser2021taming . The proposed frequency loss, $\mathcal{L}_{\text{fre}}$, is designed to capture discrepancies in the frequency domain by first transforming the input images from the spatial domain using the Fast Fourier Transform (FFT). The transformation process can be expressed as:
$\hat{I}_{\text{in}}=\mathcal{F}(I_{\text{in}}),\hat{I}_{\text{rec}}=\mathcal{F}(I_{\text{rec}}),\mathcal{F}(u,v)=\sum_{x=0}^{H-1}\sum_{y=0}^{W-1}I(x,y)\cdot e^{-i\frac{2\pi ux}{H}}\cdot e^{-i\frac{2\pi vy}{W}}.$ (15)
Next, we shift the zero-frequency components of both $\hat{I}_{\text{in}}$ and $\hat{I}_{\text{rec}}$ which represent the average intensity of the images to the center of the spectrum, resulting in $\hat{I}_{\text{in}}$ and $\hat{I}_{\text{rec}}$ for loss computation. The loss can then be formulated as:
$\mathcal{L}_{\text{fre}}=\left(N\left(\log(1+|\hat{I}^{\text{shift}}_{\text{in}}|)\right)-N\left(\log(1+|\hat{I}^{\text{shift}}_{\text{rec}}|)\right)\right)^{2},\$(16)
where $N(\cdot)$ denotes a normalization operation. Optimizing the above losses encourages the VAE to focus on semantic information relevant to fusion during image reconstruction.
VAE training stage II To address Challenge II, we propose a training strategy for Variational Autoencoder (VAE) tailored to the image fusion task. Specifically, we integrate the VAE into the overall fusion framework and perform joint training to enhance its adaptability to the fusion process. It is important to note that, in our method, the input is a visible image, and prior information from the fused image is incorporated during the sampling stage to achieve the fusion. As a result, the VAE encoder is only required to effectively compress the input image, while the decoder is responsible for both reconstructing the image and incorporating fusion-related information. Therefore, in the second stage of training, we focus on fine-tuning the VAE decoder to improve its ability to reconstruct fused images. During this stage, we employ a fusion-specific loss function commonly used in image fusion tasks to optimize the VAE, which is formulated as follows:
$\mathcal{L}_{\text{fusion}}=\lambda_{\text{int}}\mathcal{L}_{\text{int}}+\lambda_{\text{SSIM}}\mathcal{L}_{\text{SSIM}}+\lambda_{\text{grad}}\mathcal{L}_{\text{grad}}+\lambda_{\text{color}}\mathcal{L}_{\text{color}}+\lambda_{\text{mask}}\mathcal{L}_{\text{mask}},$ (17)
where $\mathcal{L}_{\text{int}}$, $\mathcal{L}_{\text{SSIM}}$, $\mathcal{L}_{\text{grad}}$, and $\mathcal{L}_{\text{color}}$ are followed by yi2024text . Meanwhile, to achieve saliency-guided regional fusion, we introduce a saliency mask loss, denoted as $\mathcal{L}_{\text{mask}}$, which can be formulated as:
$\mathcal{L}_{\text{mask}}=\left\|\mathcal{W}_{v}\cdot I_{v}+\mathcal{W}_{ir}\cdot I_{ir}-I_{f}\right\|_{1}.$ (18)
Among them, $I_{ir}$, $I_{v}$ represent the input images, and $I_{f}$ are the fused image. $\mathcal{W}_{v}$ and $\mathcal{W}_{ir}$ denote the saliency-based weight maps computed from the corresponding input images. The saliency mask loss $\mathcal{L}_{\text{mask}}$ guides the network to focus on salient regions during the fusion process, thereby improving the preservation of complementary information within the fused image. By jointly optimizing this loss with other fusion objectives, the reconstruction capability of the VAE decoder is significantly enhanced, ultimately leading to high-quality image fusion results. More details about the training loss $\mathcal{L}_{\text{fusion}}$ and $\mathcal{L}_{\text{mask}}$ can be found in Appendix B.
Figure 3: Visual comparison of IVIF with SOTA methods on M3FD datasets.
5 Experiments
Table 1: Comparison of Metrics with Our Baseline Method DDFM zhao2023ddfm .
  Methods M3FD Dataset T&R Datasets
EN MI SF VIF SSIM EN MI SF VIF SSIM
DDFM 6.720 2.871 9.102 0.677 0.867 7.077 1.798 8.910 0.277 0.207
Ours 6.722 3.320 9.780 0.748 0.914 7.139 2.948 12.55 0.675 0.921
(+0.002) (+0.449) (+0.678) (+0.071) (+0.047) (+0.062) (+1.150) (+3.640) (+0.398) (+0.714)
Experiment datasets We conduct experiments on three representative image fusion tasks: infrared and visible image fusion (IVIF), multi-exposure image fusion (MEF), and multi-focus image fusion (MFF). For the infrared and visible image fusion task, evaluations are performed on three widely-used benchmark datasets: $\text{M}^{3}\text{FD}$ liu2022target , TNO toet2012progress , and RoadScene xu2020fusiondn . For the multi-exposure and multi-focus fusion tasks, we utilize the MEFB zhang2021benchmarking and MFIF zhang2021deep datasets, respectively. The MFIF dataset includes the Lytro nejati2015multi , MFFW xu2020mffw , and MFI-WHU zhang2021mff datasets.
Implementation Details The two-stage training of the VAE was conducted entirely on an NVIDIA V100 GPU. In the first stage, the model was trained on the LLVIP jia2021llvip and MSRS tang2022piafusion datasets for 20 epochs. Interestingly, the best validation performance was typically achieved within just 4 to 5 epochs. The second stage involved training exclusively on the MSRS tang2022piafusion dataset for 40 epochs. The remaining hyperparameters for both stages were configured in accordance with the experimental settings detailed in rombach2022high and yi2024text . We evaluate our method on all three fusion tasks using the same set of checkpoints, without any task-specific fine-tuning, thereby demonstrating the strong generalization capability of our approach across diverse tasks.
Table 2: Quantitative comparison on M3FD, TNO, and RoadScene datasets. The best and second best results are highlighted in bold and underline.
     Dataset M3FD Dataset       T&R Dataset
Method MI $\uparrow$ VIF $\uparrow$ SCD $\uparrow$ EN $\uparrow$ MI $\uparrow$ VIF $\uparrow$ SCD $\uparrow$ EN $\uparrow$
U2Fusion xu2020u2fusion 2.760 0.633 1.569 6.659 2.599 0.556 1.338 6.821
YDTR tang2022ydtr 3.183 0.635 1.506 6.547 2.976 0.588 1.420 6.842
UMFusion wang2022unsupervised 3.089 0.613 1.570 6.669 2.888 0.610 1.475 6.967
ReCoNet huang2022reconet 3.066 0.577 1.483 6.679 2.985 0.540 1.510 7.051
LRRNet li2023lrrnet 2.805 0.566 1.463 6.437 2.766 0.508 1.558 7.118
CoCoNet liu2024coconet 2.631 0.729 1.772 7.738 2.579 0.568 1.782 7.735
DDFM zhao2023ddfm 2.871 0.677 1.683 6.720 1.798 0.277 1.160 7.077
Ours 3.320 0.748 1.574 6.722 2.948 0.675 1.639 7.139
Table 3: Efficiency comparisons with other diffusion-based methods. The best and second best results are highlighted in bold and underline.
  Metrics Methods
DRMF Dif-Fusion Diff-IF Text-DiFuse DDFM CCF Ours
[…]
Runtime (s) $\downarrow$ 3.221 1.997 2.457 9.199 22.03 62.47 0.308
Parameters (M) 170.98 416.47 23.47 119.49 552.81 552.81 65.57
5.1 Experiments on Infrared and Visible Image Fusion
In this section, we conduct a comprehensive comparison between our proposed RFfusion method and other fusion approaches. We begin by comparing it with our baseline method, DDFM zhao2023ddfm . Subsequently, we evaluate its performance against several state-of-the-art methods proposed in recent years to demonstrate the superiority of our approach, including: U2Fusion xu2020u2fusion , YDTR tang2022ydtr , UMFusion wang2022unsupervised , ReCoNet huang2022reconet , LRRNet li2023lrrnet , CoCoNet liu2024coconet , and DDFM zhao2023ddfm .
Comparison with DDFM method Since our method is built upon the DDFM framework by introducing fusion priors to achieve image fusion—with the main differences lying in the sampling strategy and the use of VAE for latent space generation—we primarily compare our approach with DDFM. As shown in Table 1, our method significantly accelerates inference and reduces computational overhead, while outperforming DDFM across all fusion metrics on multiple datasets. These results indicate that our method not only effectively reduces the number of sampling steps but also substantially enhances the quality of the fused images. Furthermore, the results validate the generality and flexibility of our approach, demonstrating its potential to serve as a plug-and-play module that can be integrated into other diffusion-based image fusion frameworks to simultaneously improve both inference efficiency and fusion performance.
Quantitative Comparison As shown in Table 2, we conducted a comprehensive evaluation of the proposed method using four widely adopted quantitative metrics across three benchmark datasets: M3FD, TNO, and Roadscene. The results demonstrate that our method consistently ranks among the top two across most metrics, indicating strong overall performance. Specifically, on the M3FD dataset, our method achieves the best performance in terms of Mutual Information (MI), highlighting its effectiveness in preserving informative content from the source images. Moreover, our method achieves the highest scores in Visual Information Fidelity (VIF) across all datasets, further confirming its superiority in enhancing the visual quality of the fused images.
Qualitative Comparison As shown in Figure 3, our method demonstrates superior visual performance compared to other approaches. We selected four images from the M3FD dataset for qualitative analysis, covering different scenarios including both daytime and nighttime, to ensure a comprehensive evaluation. Our method better preserves detailed texture information from the original images, such as window textures on buildings and fine details at the ends of tree branches, whereas other methods tend to blur these features. Additionally, our method highlights human details more effectively and better retains the mutual information between visible and infrared images. This demonstrates the advantages of our method in qualitative results.
Figure 4: Visual comparison of multi-exposure image fusion and multi-focus image fusion with SOTA methods on MEFB and MFIF datasets.
5.2 Experiments on efficiency comparisons with other diffusion-based methods
To verify the effectiveness of our method in reducing the inference time and computational cost of diffusion models in image fusion tasks, we compare it with several diffusion-based image fusion approaches proposed in recent years, including DRMF tang2024drmf , Dif-Fusion yue2023dif , Diff-IF yi2024diff , Text-DiFuse zhang2024text , DDFM zhao2023ddfm , and CCF cao2024conditional . All experiments are conducted on an NVIDIA V100 GPU, and the fusion speed as well as the number of model parameters are evaluated on the RoadScene xu2020fusiondn dataset to comprehensively assess the efficiency and complexity of each method. As shown in Table 3 and Figure 1, compared with other diffusion-based image fusion methods, our approach demonstrates a significant advantage in inference speed while also achieving superior fusion quality. These results indicate that our method not only greatly improves inference efficiency but also maintains, or even enhances fusion performance, fully validating its capability for joint optimization of efficiency and effectiveness.
Table 4: Quantitative comparison on MEFB and MFIF datasets. The best and second best results are highlighted in bold and underline.
     Dataset MEFB Dataset       MFIF Dataset
Method MI $\uparrow$ $C$ $\uparrow$ Qcb $\uparrow$ PSNR $\uparrow$ MI $\uparrow$ $C$ $\uparrow$ Qcb $\uparrow$ PSNR $\uparrow$
DeFusion liang2022fusion 4.854 0.834 0.365 57.70 6.007 0.976 0.627 76.62
TC-MoA zhu2024task 5.418 0.900 0.430 59.00 6.686 0.968 0.731 74.78
Text-IF yi2024text 5.596 0.860 0.385 56.44 5.399 0.967 0.629 71.74
DDFM zhao2023ddfm 3.850 0.792 0.321 58.39 3.232 0.772 0.413 66.24
CCF cao2024conditional 4.830 0.898 0.398 58.38 4.799 0.956 0.474 66.64
Ours 6.528 0.901 0.461 58.49 6.443 0.977 0.654 75.04
5.3 Experiments on Evaluation on Multi-Focus Fusion
Quantitative Comparison As shown in Table 4, we conducted a comprehensive evaluation of the proposed method on the MFIF dataset. The experimental results demonstrate that our method consistently ranks among the top two across most evaluation metrics, fully validating its superior effectiveness in the multi-focus fusion (MFF) task. Notably, the proposed method achieves this performance without any fine-tuning on the multi-focus fusion dataset, further confirming its strong generalization ability and robustness across different tasks.
Qualitative Comparison As shown in Figure 4, we selected two representative images from the MFIF dataset for qualitative analysis, covering both daytime and nighttime scenarios to ensure a comprehensive evaluation. Compared to other methods, our method more effectively preserves clear details from the original images, achieving high-quality multi-focus image fusion and fully demonstrating its superior performance in qualitative evaluation.
5.4 Experiments on Evaluation on Multi-Exposure Fusion
Quantitative Comparison Table 4 presents a quantitative comparison between our method and existing approaches on the MEFB dataset. Notably, our method does not require any fine-tuning on the MEF dataset. It consistently outperforms other multi-exposure fusion (MEF) methods across most evaluation metrics, demonstrating its superior performance and strong generalization capability in the multi-exposure image fusion task.
Qualitative Comparison As shown in Figure 4, compared to other methods, our fusion results better preserve the detailed features of the original images and achieve superior visual performance. We selected two images from the MEFB dataset for qualitative analysis. Our method more effectively retains the texture of windows and the contour features of candles, demonstrating its advantages in qualitative evaluation.
Table 5: Ablation studies on the effectiveness of the two-stage training strategy and loss functions.
  Stage I Stage II PSNR MI SF AG
59.41 2.998 12.16 4.615
59.68 3.017 12.88 4.676
60.36 3.001 12.57 4.783
61.81 3.220 14.00 5.218
  $\mathcal{L}_{\text{fre}}$ $\mathcal{L}_{\text{mask}}$ PSNR MI SF AG
57.22 2.944 12.77 4.882
58.84 3.202 13.56 5.021
59.67 3.121 13.31 4.976
61.81 3.220 14.00 5.218
5.5 Ablation Study
Experimental on the effectiveness of the two-stage training strategy. We conducted ablation studies on the two-stage training strategy for the VAE to evaluate its effectiveness in the image fusion task. As shown in Table 5, we compared fusion performance under four settings: without any training, using only the first-stage training, using only the second-stage training, and applying both stages of training. The results demonstrate that the fusion performance is optimal when both stages are applied, validating the effectiveness of the training strategy in enhancing fusion quality.
Experimental on the effectiveness of the loss functions. We conducted ablation experiments on the loss functions used in our proposed training method to evaluate their contributions to image fusion performance, as shown in Table 5. Specifically, we designed four experimental settings: without using either $\mathcal{L}_{\text{fre}}$ or $\mathcal{L}_{\text{mask}}$, using only $\mathcal{L}_{\text{fre}}$, using only $\mathcal{L}_{\text{mask}}$, and using both $\mathcal{L}_{\text{fre}}$ and $\mathcal{L}_{\text{mask}}$ simultaneously. The experimental results demonstrate that both $\mathcal{L}_{\text{fre}}$ and $\mathcal{L}_{\text{mask}}$ can independently improve fusion quality, while their combined use leads to the best performance. These findings validate the effectiveness and necessity of the proposed loss function design.
6 Limitation
RFfusion still relies on a Rectified Flow pre-trained model trained on generic image generation tasks, which is not specifically designed for image fusion. This limitation may hinder the further improvement of RFfusion’s fusion performance.
7 Conclusion
In this paper, we propose an efficient one-step diffusion-based image fusion method called RFfusion. By integrating Rectified Flow into the image fusion task, our method leverages its efficient one-step sampling mechanism to significantly accelerate the diffusion-based fusion process. Moreover, we design a task-specific Variational Autoencoder (VAE) that performs fusion in the latent space, effectively reducing computational overhead while preserving more image details. Extensive experimental results demonstrate that RFfusion achieves superior performance in both inference speed and fusion quality compared to existing state-of-the-art methods, and also exhibits strong generalization capabilities across diverse image fusion tasks. In the future, we will further explore acceleration mechanisms of diffusion models in image fusion tasks to achieve more efficient image fusion methods.